12 liters of 8% acid solution with a 20% acid solution, which results in a 16% acid solution. find number of liters of 20% acid solution in the new mixture.
Answer by stanbon(75887)

12 liters of 8% acid solution with a 20% acid solution, which results in a 16% acid solution. find number of liters of 20% acid solution in the new mixture.
------
Equation:
acid + acid = acid
0.08*12 + 0.20x = 0.16(12+x)
--------
8*12 + 20x = 16*12 + 16x
----
4x = 8*12
x = 24 liters (amt. of 20% solution needed)
==================================
Cheers,
Stan H.